if x varies directly as y, and x=42 when y = 6, find x when y = 7: if x varies directly as y, and x=42 when y = 6, find x when y = 7
Answer by checkley77(3397)

42/6=x/7
6x=7*42
6x=294
x=294/6
x=49 answer.